OpenStudy (anonymous):

Please help Medal! For the system shown below what are the coordinates of the solution that lies in Quadrant 2? Write your answer in the form (a,b) x^2+4y^2=100 4y^2-x^2=-20

OpenStudy (anonymous):

\[8y^2 = 80; y^2 = 10; y =\+/-sqrt(10)\] plug in 10 for y^2 \[x^2+4(10) = 100; x^2 = 60; x = +/- \sqrt(60)\] In quadrant II, y is positive and x is negative So the coordinates are \[(-\sqrt(60), \sqrt(10))\]

OpenStudy (anonymous):

Do you know how I got \[8y^2 = 80\]?

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

Add the two equations together. \[4y^2+4y^2 = 8y^2\] \[x^2-x^2 = 0\] \[100+ (-20) = 80\] Do you see?

OpenStudy (anonymous):

I just added like terms.

OpenStudy (anonymous):

ok yea i see it

OpenStudy (anonymous):

wait omg i wrote one of the equations wrong!

OpenStudy (anonymous):

my mistake this is the correct equations x^2+4y^2=100 4y-x^2-20

OpenStudy (anonymous):

@DoYourHomework

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Wait. Is it 4y-x^2 = -20?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

my keyboard isnt working properly

OpenStudy (anonymous):

Add the two equations to get. \[4y^2+4y = 80\] Divide equation by 4 \[y^2+y = 20\] Bring 20 on other side and solve for y. \[y^2+y-20 = (y+5)(y-4); y = -5 and y = 4 \] Since we are looking at the second quadrant, y must be positive. So, y = 4 Plug y into either equation \[4(4)-x^2 = -20; -x^2 = -36; x^2 = 36; x = +/-6\] Since x is negative in quadrant 2 the coordinates are (-6,4)

OpenStudy (anonymous):

woa thats great

OpenStudy (anonymous):

idk how you could solve it quick

OpenStudy (anonymous):

Idk, I've seen a lot of these problems before. Just apply what you learned in class that day to solve your homework problems. It's dumb advice, but it's true.

OpenStudy (anonymous):

*your homework problems for that day.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

well thank you

OpenStudy (anonymous):

np!